# Roberto Bruno, Roberto De Prisco, and Ugo Vaccaro

### v2

@@ I Introduction
Among search procedures, the binary search stands out as an important algorithmic technique to solve the computational problem whereby an unknown element $x$ is located within a finite search space $S$, and the objective is to identify $x$ by performing queries that yield Yes/No answers. These kind of queries typically fall into one of two categories: comparison queries or membership queries. In the case of comparison queries, the unknown element $x$ resides in an ordered set that can be represented as $[n]=\{1,\ldots,n\}$ The queries then take the form: ”Is the unknown number $x$ greater than $j$, where $j\in[n]$?” (see [30] for an encyclopedic treatment of the subject). For membership queries, the search space $S$ may be unordered. The queries are phrased as: ”Does the unknown number $x$ belong to a subset $A\subseteq S$?” This latter problem has also been extensively studied (see, for example, [1, 2, 33]) and has significant connections to a variety of fields, including Exact Learning [10], Combinatorics [26], and Group Testing [13], to name a few. It is well known that binary search procedures that proceed with comparison queries are mathematically equivalent to binary alphabetic codes, and that binary search procedures that proceed with membership queries are equivalent to binary prefix codes (see, e.g., [1]).
@@
    the leaves of the tree, read from the leftmost to the rightmost leaf, are in one-to-one correspondence with the elements of the ordered set $[n]=\{1,\ldots,n\}$.
@@
We remark that problems of the kind described above arise in many other different scenarios (e.g., [34, 37, 38, 39, 40, 42, 43, 44, 45, 49]) where different test outcomes might have different costs. In our case, a Yes outcome costs 0, while a No has some cost $>0$ (like in the classic egg-dropping puzzle [28], for example), and one seeks average-case optimal algorithms (with respect to the number of performed tests) with a fixed bound on the worst-case cost of the algorithm. Related problems are also studied in [20, 25] and references therein quoted.
@@ I-A Our results
The interest in the above result is twofold. First, it represents a natural extension of the classical Kraft inequality (see, eg., Thm. 5.1 of [1]) that establishes a necessary and sufficient condition for the existence of an $n$-leaf full binary tree with multiset of root-to-leaf path lengths equal to $\{\ell_{1},\ldots,\ell_{n}\}$. Secondly, the above result will be instrumental to the derivation of an improvement of the result by Dolev et al. [12] that we have discussed before. More precisely, in Section V we design a novel algorithm to construct a minimum average height tree that satisfies properties 1’, 2’, and 3’; our algorithm has $O(n^{2}D)$ time complexity, considerably improving on the $O(n^{2+D})$ algorithm presented in [12]. Note that the parameter $D$ can possibly assume any value in the interval $[1,\log n]$; therefore, our algorithm is the first polynomial in $n$ time algorithm to solve the problem discussed before.
@@ II Preliminaries
    no codeword $w(s)$ is prefix of another $w(s^{\prime})$, for any $s,s^{\prime}\in[n]$, $s\neq s^{\prime}$.
@@
[…]
@@ III A Dynamic Programming approach
We remind that the right subtree is reached from a right-pointing edge of the root; therefore, all the corresponding codewords one obtains begin with symbol 1, and this accounts for the value $w-1$ in the parameters of the right subtree cost.
@@ III-A An improved algorithm
$F(i,j,r)=\begin{cases}w(i),&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ if }i=j\text{ and }r\geq 0,\\ \min_{i\leq k<j}\{aF(i,k,f(r))+bF(k+1,j,g(r))\\ \qquad+h(i,k,j)\},&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ if }f(r)\geq 0\text{ and }g(r)\geq 0,\\ \mbox{undefined},&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ otherwise,}\end{cases}$ (4)
@@
We need to show that the function $C$ is a special case of (4). For such a purpose, let $w(i)=0$ for each $i=1,\dots,n$, and let $a=b=1$. While as functions $f$ and $g$ we take $f(x)=x$ and $g(x)=x-1$. Finally, we define the function $h(i,k,j)=\text{c}(i,j)$ for each $i\leq k<j$, where $\text{c}(i,j)=p_{i}+\dots+p_{j}$. One can see that the function $h$ defined in this way satisfies the conditions (5-8). In fact, for $i\leq j\leq t<s\leq l$ and $i\leq s$, if $t\leq k$ it holds that
@@
This is the quadrangle inequality for the function $C$ with parameter $w-1$, with $k\leq k^{\prime}\leq j<j+1$, and from Corollary 1 we know that it holds true. Thus, this concludes the proof. ∎
Due to Theorem 1, instead of searching from $i+1$ to $j$ (line 18 of Algorithm 1), it is sufficient to search from $R(i,j-1,w)$ through $R(i+1,j,w)$. Therefore line 18 of Algorithm 1 can be substituted by “for $k\leftarrow R(i,j-1,w)$ to $R(i+1,j,w)$”. Exactly as argued by Knuth (see the remark after the Corollary, p. 19 of [29]), when $C(i,j,w)$ is being calculated with this modification, only $R(i+1,j,w)-R(i,j-1,w)+1$ cases need to be examined. Summing for fixed $j-i$ gives a telescoping series showing that the total amount of work of each iteration of the loop at line 12 of the modified Algorithm 1 is proportional to $O(n^{2})$, at worst. This decreases the total time complexity of the algorithm from $O(n^{3}D)$ to $O(n^{2}D)$.
@@ IV A Kraft-like condition for the existence of constrained binary trees
Let $\{\ell_{1},\ldots,\ell_{n}\}$ be a multiset of positive integers, $D$ a positive integer and $L=\max_{i}\ell_{i}$. Let $N_{j}$ count the number of integers in $\{\ell_{1},\ldots,\ell_{n}\}$ that are equal to $j$, for $j=1,\dots,L$. Then, there exists a binary prefix code with codeword lengths $\ell_{1},\ldots,\ell_{n}$, such that each codeword does not contain more than $D$ $1$’s, if the following inequalities hold:
@@
We label each left-pointing edge of $T$ with 00 and each right-pointing edge with $1$. The set of all binary words one gets by reading the sequences of binary labels from the root of $T$ to each leaf gives a binary prefix code $P$ in which ${\lvert w\rvert_{1}}\leq D$, for each $w\in P.$ We now show that the number of nodes at the level $j$ of such a tree $T$ is exactly equal to
@@
First of all, we recall that each path in the tree can be seen as a binary string. Therefore, (25) holds because the first term $\sum_{i=0}^{D-1}\binom{j}{i}$ counts the number of binary strings of length $j$, each containing at most $D-1$ ones. The number of binary strings of length $j$ containing exactly $D$ ones that one obtains, by reading the root-to-leaves path in $T$ is equal to $\binom{j-1}{D-1}$. This is true because any binary string of length $j$ that one obtains by reading the root-to-leaves path in $T$ and that contains $D$ ones must end with 1. In fact, let us consider a binary string of length $j$ that contains exactly $D$ ones and does not end with 1. In the path from the root to the node associated with such a binary string, there will be at least one internal node that has only exactly one child. Therefore, such a node cannot belong at level $j$ of the binary tree $T$ since, by hypothesis, $T$ is full.
@@
Hence, there are enough available nodes in the level $L$ of $T$ to construct $N_{L}$ codewords of length $L$. In particular, it will be convenient (for future purposes) to construct such codewords by choosing the first $N_{L}$ nodes, in sequence, starting from the left-most one. The constraint that the obtained codewords $w$’s have to satisfy (i.e., that ${\lvert w\rvert_{1}}\leq D$) is automatically satisfied since in the tree $T$ any root-to-leaf path has at most $D$ right-pointing edges, each labeled with $1$.
The nodes at level $L-1$ that are parents of the chosen nodes at level $L$ cannot be used anymore in the construction of the code, to comply with the prefix constraint of $C$. Therefore, we need to ignore such nodes at level $L-1$ and their number is exactly $M_{L}=\lceil N_{L}/2\rceil.$ Thus, the number of nodes at level $L-1$ that can be still used is
@@
therefore we can iterate the same process to construct the desired number of codewords of length $L-1$. We choose the first $N_{L-1}$ available nodes at level $L-1$ (that are all the nodes at level $L-1$ that give rises to not-prefixes of the chosen nodes at level $L$) in sequence starting from the left-most one. Again, we have to ignore all the nodes at level $L-2$ that are ancestors of the chosen nodes at level $L$ and $L-1$. This number is exactly
@@
Since the inequalities hold, we can construct the prefix code with the required list of lengths $\langle 4,2,3,4,3,3,3\rangle$. For such a purpose, let us start from the complete binary tree $T$ of depth $4$ in which, for any node $v$, the path from the root to the node $v$ does not contain more than $D=2$ edge with label $1$ (Figure 1).
Figure 1: The complete binary tree $T$ of depth $4$
Since $N_{4}=2$ we need two nodes at the level $4$ of $T$ to construct the codewords for the symbols $s_{1}$ and $s_{4}$, respectively. As described in Lemma 2, we take the left-most ones and we get the tree in Figure 2. Note that is the inequality (31) that assures us that we can choose such nodes at the level $4$.
@@
Now, since $N_{3}=4$ we need to take four nodes at level $3$ for the construction of the codewords for the symbols $s_{3},s_{5},s_{6}$ and $s_{7}$. However, we have to ignore the nodes that are prefixes of the chosen codewords of length $4$. Note that we can do this since the values $N_{i}$’s satisfy the required inequalities.
@@
Figure 3: The tree after choosing the nodes for the codewords of length $4$ and $3$
Since $N_{2}=1$ there is only one codeword of length 2 that we need to choose. Therefore, we take the left-most node at level $2$ that is still “available” as shown in Figure 4.
@@
Figure 4: The tree after choosing the nodes for the codewords of length $4,3$ and $2$
Finally, since $N_{1}=0$ we do not need to take any node at level $1$. So, the tree that we get in the end is shown in Figure 5.
@@
is the maximum number of nodes that a complete binary tree (in which any root-to-leaf path has at most $D$ right-pointing edges) can have at level $j=1,\dots,L$. Moreover, since we are considering only full binary trees, one can see that $M_{j}$, for each $j=2,\dots,L$, corresponds to the minimum number of nodes at level $j-1$ that are necessarily prefixes of some nodes associated with codewords of length $\geq j$, regardless of how such nodes are chosen. Putting all together, since by hypothesis, $\ell_{1},\dots,\ell_{n}$ are the codeword lengths of a prefix code $C$ in which each codeword does not contain more than $D$ ones and whose binary tree representation is full, we have that the number of leaves at level $j$ in the binary tree of $C$, which is $N_{j}$, cannot be greater than
@@
for each $j=1,\dots,L$. This holds because (37) is obtained by subtracting from the total number of nodes at level $j$ (counted by (36)) the minimum number of nodes at level $j-1$ that are necessarily prefixes of some nodes associated with codewords of length $\geq j$ (counted by $M_{j+1}$). In other words, (37) represents the maximum number of leaves at level $j=1,\dots,L$, that a complete binary tree — where any root-to-leaf path has at most $D$ right-pointing edges — has still “available” (i.e., they are not prefixes of the nodes associated with codewords of length $\geq j$). ∎
@@ V Optimal Average-Length Constrained Trees
[…]
@@ V-A When the constraint is really a constraint?
The constraint that at most $D$ ones must appear in each codeword might prevent the construction of the (otherwise unrestricted) optimal tree. However, if $D$ is large enough, the best tree satisfying the constraint can be the same as the optimal (unrestricted) one. For the case of alphabetic trees establishing the smallest value of $D$ that does not prevent the construction of the optimal tree is straightforward. Indeed, it is easy to identify probability distributions (for example, a decreasing dyadic distribution) for which any optimal alphabetic code must have a codeword with $n-1$ ones. Hence, any value of $D$ strictly less than $n-1$ is a constraint that prevents the construction of a tree with the same cost of an optimal unrestricted one.

### v3

@@ I Introduction
Among search procedures, the binary search stands out as an important algorithmic technique to solve the computational problem whereby an unknown element $x$ is located within a finite search space $S$, and the objective is to identify $x$ by performing queries that yield Yes/No answers. These kind of queries typically fall into one of two categories: comparison queries or membership queries. In the case of comparison queries, the unknown element $x$ resides in an ordered set that can be represented as $n]=\{1,\ldots,n\}$ The queries then take the form: ”Is the unknown number $x$ greater than $j$, where $j\in[n]$?” (see [30] for an encyclopedic treatment of the subject). For membership queries, the search space $S$ may be unordered. The queries are phrased as: ”Does the unknown number $x$ belong to a subset $A\subseteq S$?” This latter problem has also been extensively studied (see, for example, [1, 2, 33]) and has significant connections to a variety of fields, including Exact Learning [10], Combinatorics [26], and Group Testing [13], to name a few. It is well known that binary search procedures that proceed with comparison queries are mathematically equivalent to binary alphabetic codes, and that binary search procedures that proceed with membership queries are equivalent to binary prefix codes (see, e.g., [1]).
@@
    the leaves of the tree, read from the leftmost to the rightmost leaf, are in one-to-one correspondence with the elements of the ordered set $n]=\{1,\ldots,n\}$.
@@
We remark that problems of the kind described above arise in many other different scenarios (e.g., [34, 37, 38, 39, 40, 42, 43, 44, 45, 49]) where different test outcomes might have different costs. In our case, a Yes outcome costs 0, while a No has some cost >0>0 (like in the classic egg-dropping puzzle [28], for example), and one seeks average-case optimal algorithms (with respect to the number of performed tests) with a fixed bound on the worst-case cost of the algorithm. Related problems are also studied in [20, 25] and references therein quoted.
@@ I-A Our results
The interest in the above result is twofold. First, it represents a natural extension of the classical Kraft inequality (see, eg., Thm. 5.1 of [1]) that establishes a necessary and sufficient condition for the existence of an $n$-leaf full binary tree with multiset of root-to-leaf path lengths equal to $\{\ell_{1},\ldots,\ell_{n}\}$. Secondly, the above result will be instrumental to the derivation of an improvement of the result by Dolev et al. [12] that we have discussed before. More precisely, in Section V we design a novel algorithm to construct a minimum average height tree that satisfies properties 1’, 2’, and 3’; our algorithm has $O(n^{2}D)$ time complexity, considerably improving on the $O(n^{2+D})$ algorithm presented in [12]. Note that the parameter $D$ can possibly assume any value in the interval $1,\log n]$; therefore, our algorithm is the first polynomial in $n$ time algorithm to solve the problem discussed before.
@@ II Preliminaries
    no codeword $w(s)$ is prefix of another $w(s^{\prime})$, for any $^{\prime}\in[n], sss\neq s^{\prime}$.
@@
$\mbox{\ \ \ \ subj. to }{\lvert w(s_{i})\rvert_{1}}\leq D,\quad\forall i=1,\dots,n,$
@@ III A Dynamic Programming approach
We remind that the right subtree is reached from a right-pointing edge of the root; therefore, all the corresponding codewords one obtains begin with symbol 1, and this accounts for the value w1w-1 in the parameters of the right subtree cost.
@@ III-A An improved algorithm
[…]
@@
We need to show that the function $C$ is a special case of (4). For such a purpose, let $w(i)=0$ for each $i=1,\dots,n$, and let $a=b=1$. While as functions $f$ and $g$ we take $f(x)=x$ and g(x)=x1g(x)=x-1. Finally, we define the function $h(i,k,j)=\text{c}(i,j)$ for each $i\leq k<j$, where $\text{c}(i,j)=p_{i}+\dots+p_{j}$. One can see that the function $h$ defined in this way satisfies the conditions (5-8). In fact, for $i\leq j\leq t<s\leq l$ and $i\leq s$, if $t\leq k$ it holds that
@@
This is the quadrangle inequality for the function $C$ with parameter w1w-1, with $k\leq k^{\prime}\leq j<j+1$, and from Corollary 1 we know that it holds true. Thus, this concludes the proof. ∎
Due to Theorem 1, instead of searching from $i+1$ to $j$ (line 18 of Algorithm 1), it is sufficient to search from R(i,j1,w)R(i,j-1,w) through $R(i+1,j,w)$. Therefore line 18 of Algorithm 1 can be substituted by “for $k\leftarrow R(i,j-1,w)$ to R(i+1,j,w)R(i+1,j,w)”. Exactly as argued by Knuth (see the remark after the Corollary, p. 19 of [29]), when $C(i,j,w)$ is being calculated with this modification, only R(i+1,j,w)R(i,j1,w)+1R(i+1,j,w)-R(i,j-1,w)+1 cases need to be examined. Summing for fixed jij-i gives a telescoping series showing that the total amount of work of each iteration of the loop at line 12 of the modified Algorithm 1 is proportional to $O(n^{2})$, at worst. This decreases the total time complexity of the algorithm from $O(n^{3}D)$ to $O(n^{2}D)$.
@@ IV A Kraft-like condition for the existence of constrained binary trees
Let $\{\ell_{1},\ldots,\ell_{n}\}$ be a multiset of positive integers, $D$ a positive integer and $L=\max_{i}\ell_{i}$. Let $N_{j}$ count the number of integers in $\{\ell_{1},\ldots,\ell_{n}\}$ that are equal to $j$, for $j=1,\dots,L$. Then, there exists a binary prefix code with codeword lengths $\ell_{1},\ldots,\ell_{n}$, such that each codeword does not contain more than $D$ 11’s, if the following inequalities hold:
@@
We label each left-pointing edge of $T$ with 0 and each right-pointing edge with 11. The set of all binary words one gets by reading the sequences of binary labels from the root of $T$ to each leaf gives a binary prefix code $P$ in which ${\lvert w\rvert_{1}}\leq D$, for each $w\in P.$ We now show that the number of nodes at the level $j$ of such a tree $T$ is exactly equal to
@@
First of all, we recall that each path in the tree can be seen as a binary string. Therefore, (25) holds because the first term $\sum_{i=0}^{D-1}\binom{j}{i}$ counts the number of binary strings of length $j$, each containing at most D1D-1 ones. The number of binary strings of length $j$ containing exactly $D$ ones that one obtains, by reading the root-to-leaves path in $T$ is equal to $\binom{j-1}{D-1}$. This is true because any binary string of length $j$ that one obtains by reading the root-to-leaves path in $T$ and that contains $D$ ones must end with 1. In fact, let us consider a binary string of length $j$ that contains exactly $D$ ones and does not end with 1. In the path from the root to the node associated with such a binary string, there will be at least one internal node that has only exactly one child. Therefore, such a node cannot belong at level $j$ of the binary tree $T$ since, by hypothesis, $T$ is full.
@@
Hence, there are enough available nodes in the level $L$ of $T$ to construct $N_{L}$ codewords of length $L$. In particular, it will be convenient (for future purposes) to construct such codewords by choosing the first $N_{L}$ nodes, in sequence, starting from the left-most one. The constraint that the obtained codewords $w$’s have to satisfy (i.e., that ${\lvert w\rvert_{1}}\leq D$) is automatically satisfied since in the tree $T$ any root-to-leaf path has at most $D$ right-pointing edges, each labeled with 11.
The nodes at level L1L-1 that are parents of the chosen nodes at level $L$ cannot be used anymore in the construction of the code, to comply with the prefix constraint of $C$. Therefore, we need to ignore such nodes at level L1L-1 and their number is exactly $M_{L}=\lceil N_{L}/2\rceil.$ Thus, the number of nodes at level L1L-1 that can be still used is
@@
therefore we can iterate the same process to construct the desired number of codewords of length L1L-1. We choose the first $N_{L-1}$ available nodes at level L1L-1 (that are all the nodes at level L1L-1 that give rises to not-prefixes of the chosen nodes at level $L$) in sequence starting from the left-most one. Again, we have to ignore all the nodes at level L2L-2 that are ancestors of the chosen nodes at level $L$ and L1L-1. This number is exactly
@@
Since the inequalities hold, we can construct the prefix code with the required list of lengths $\langle 4,2,3,4,3,3,3\rangle$. For such a purpose, let us start from the complete binary tree $T$ of depth 44 in which, for any node $v$, the path from the root to the node $v$ does not contain more than $D=2$ edge with label 11 (Figure 1).
Figure 1: The complete binary tree $T$ of depth 44
Since $N_{4}=2$ we need two nodes at the level 44 of $T$ to construct the codewords for the symbols $s_{1}$ and $s_{4}$, respectively. As described in Lemma 2, we take the left-most ones and we get the tree in Figure 2. Note that is the inequality (31) that assures us that we can choose such nodes at the level 44.
@@
Now, since $N_{3}=4$ we need to take four nodes at level 33 for the construction of the codewords for the symbols $s_{3},s_{5},s_{6}$ and $s_{7}$. However, we have to ignore the nodes that are prefixes of the chosen codewords of length 44. Note that we can do this since the values $N_{i}$’s satisfy the required inequalities.
@@
Figure 3: The tree after choosing the nodes for the codewords of length 44 and 33
Since $N_{2}=1$ there is only one codeword of length 2 that we need to choose. Therefore, we take the left-most node at level 22 that is still “available” as shown in Figure 4.
@@
Figure 4: The tree after choosing the nodes for the codewords of length $4,3$ and 22
Finally, since $N_{1}=0$ we do not need to take any node at level 11. So, the tree that we get in the end is shown in Figure 5.
@@
is the maximum number of nodes that a complete binary tree (in which any root-to-leaf path has at most $D$ right-pointing edges) can have at level $j=1,\dots,L$. Moreover, since we are considering only full binary trees, one can see that $M_{j}$, for each $j=2,\dots,L$, corresponds to the minimum number of nodes at level j1j-1 that are necessarily prefixes of some nodes associated with codewords of length $\geq j$, regardless of how such nodes are chosen. Putting all together, since by hypothesis, $\ell_{1},\dots,\ell_{n}$ are the codeword lengths of a prefix code $C$ in which each codeword does not contain more than $D$ ones and whose binary tree representation is full, we have that the number of leaves at level $j$ in the binary tree of $C$, which is $N_{j}$, cannot be greater than
@@
for each $j=1,\dots,L$. This holds because (37) is obtained by subtracting from the total number of nodes at level $j$ (counted by (36)) the minimum number of nodes at level j1j-1 that are necessarily prefixes of some nodes associated with codewords of length $\geq j$ (counted by $M_{j+1}$). In other words, (37) represents the maximum number of leaves at level $j=1,\dots,L$, that a complete binary tree — where any root-to-leaf path has at most $D$ right-pointing edges — has still “available” (i.e., they are not prefixes of the nodes associated with codewords of length $\geq j$). ∎
@@ V Optimal Average-Length Constrained Trees
$\mbox{subj. to }{\lvert w(s_{i})\rvert_{1}}\leq D,$ $\quad\forall i=1,\dots,n,$
@@ V-A When the constraint is really a constraint?
The constraint that at most $D$ ones must appear in each codeword might prevent the construction of the (otherwise unrestricted) optimal tree. However, if $D$ is large enough, the best tree satisfying the constraint can be the same as the optimal (unrestricted) one. For the case of alphabetic trees establishing the smallest value of $D$ that does not prevent the construction of the optimal tree is straightforward. Indeed, it is easy to identify probability distributions (for example, a decreasing dyadic distribution) for which any optimal alphabetic code must have a codeword with n1n-1 ones. Hence, any value of $D$ strictly less than n1n-1 is a constraint that prevents the construction of a tree with the same cost of an optimal unrestricted one.